### v1

@@ 3.2. Task Formulation
Given a target XBRL filing with six interconnected, XML-format documents $\mathcal{D}=\{d_{1},d_{2},d_{3},d_{4},d_{5},d_{6}\}$ and a reference structured document (US-GAAP taxonomy) $\mathcal{T}$, where $d_{i}$ refers to one of the documents described in Table 2. We aim to enable LLMs to identify errors $\mathcal{E}$, including semantic matching errors, relationship extraction errors, and mathematical reasoning errors, in the target filing by understanding its internal structure and semantics in conjunction with domain knowledge derived from the reference taxonomy. We formulate it as follows:
@@
where $\mathcal{E}$ is the specific description of an error found in the $\mathcal{D}$.
To operationalize error detection across the key dimensions of XBRL filings, we identify three fundamental capabilities required for effective model evaluation: semantic consistency, structural relationship understanding, and numerical reasoning. These capabilities correspond to the primary sources of errors observed in real-world filings and capture the essential challenges in interpreting and validating structured financial data. Based on this decomposition, FinAudit defines three tasks: FinSM for identifying mismatches between textual mentions and standardized taxonomy labels, FinRE for detecting errors in hierarchical or compositional relationships among financial items, and FinMR for verifying the correctness of numerical values based on accounting logic and contextual inference. Each sub-task is associated with a specific input format tailored to its objective.
3.2.1. Financial Semantic Matching (FinSM)
The semantic matching task focuses on identifying financial tags in an XBRL filing that do not correctly align with standardized concepts defined in the US-GAAP taxonomy. The objective is to assess whether LLMs can accurately detect mismatched US-GAAP tags by filtering out semantically inconsistent or inappropriate concept assignments from a set of complex, structured documents.
This task follows the information retrieval paradigm: given a query $Q$ describing a financial term that represents either currency or concentration of credit risk, an XBRL filing $\mathcal{D}$, and a US-GAAP taxonomy $\mathcal{T}$. We formulate this task as follows:
(2) $f:(Q,\mathcal{D},\mathcal{T})\mapsto\mathcal{E}$
where $\mathcal{E}$ is the set of mismatched US-GAAP tags after retrieving.
@@ 3.2.1. Financial Semantic Matching (FinSM)
3.2.2. Financial Relationship Extraction (FinRE)
The relationship extraction task aims to identify the structural relationships among financial elements in a filing. It assesses an LLM’s ability to interpret hierarchical and compositional dependencies by aligning the document structure with the external taxonomy. This task reflects the LLMs’ capacity for semantic understanding and structured data interpretation within the context of financial reporting.
@@ 3.2.2. Financial Relationship Extraction (FinRE)
More specifically, given a relationship label space $L=\{l_{1},l_{2},l_{3}\}$ with the definitions for each relationship label, an XBRL filing $\mathcal{D}$, and a US-GAAP taxonomy $\mathcal{T}$. The goal of this task is to classify the relation type $r$ for two specific elements $e_{1}$ and $e_{2}$ based on $\mathcal{D}$ and $\mathcal{T}$. More specifically, we define three relation error types in $L$: (1) $l_{1}=Reversal$, occurs when the hierarchical relationship between sibling and child elements is mistakenly reversed; (2) $l_{2}=Inappropriateness$, indicates that a child element is incorrectly associated with an inappropriate parent element; and (3) $l_{3}=CombinationErr$, refers to an invalid combination of axis and member elements that violates the structural constraints defined in the taxonomy. Therefore, we formulate this task as follows.
(3) $f:(e_{1},e_{2},L,\mathcal{D},\mathcal{T})\mapsto\mathcal{E}$
where $\mathcal{E}$ is the specific identified relation type $r$ from the relationship label space $L$.
3.2.3. Financial Mathematical Reasoning (FinMR)
The mathematical reasoning task focuses on inferring numerical logic among financial elements based on structured representations, calculation hierarchies, and logical constraints defined in the XBRL filing. The objective is to assess whether reported values are internally consistent and compliant with domain-specific standards.
Specifically, given two questions $q_{1}$ and $q_{2}$, where $q_{1}$ concerns the extraction of a reported value and $q_{2}$ pertains to the calculation of the corresponding real value, an XBRL filing $\mathcal{D}$, and a US-GAAP taxonomy $\mathcal{T}$, the task is to extract the reported value $v$ for a given instance in $\mathcal{D}$ and to compute the numeric value $\mu$ for that instance, which is then used to verify whether the reported value $v$ is correct.
(4) $f:(q_{1},q_{2},\mathcal{D},\mathcal{T})\mapsto\mathcal{E}$
where $\mathcal{E}$ denotes the generated answer containing the extracted reported value $v$ and the computed value $\mu$, represented in JSON format.
@@ 3.2.3. Financial Mathematical Reasoning (FinMR)
4. Data Annotation and Evaluation Setting
4.1. Raw Data Collection
To create this benchmark, we collected a total of 372 XBRL filings, which were processed through two steps: pre-screening and filtering. Pre-screening serves as the foundation for filtering, enabling the identification and collection of common errors in XBRL filings.
For the pre-screening step, we accessed the XBRL US official website [^2] https://xbrl.us/data-quality/filing-results/ and randomly collected 4,545 error messages from 372 companies spanning the years 2020 to 2024. These messages, linked to the DQC US rules [^3] https://github.com/DataQualityCommittee/dqc_us_rules/tree/master, were analyzed to assess rule coverage and to identify the most common error types. Our analysis revealed that the 4,545 messages contain more than 100 distinct error types. However, as shown in Table 3, 9 error types occur most frequently, accounting for approximately 60.33% of all cases. These errors can be grouped into three broad categories: semantic-based, relation-based, and calculation-based.
For the filtering step, we extracted 1,102 error messages from the initial set of 4,545 as the first version of our benchmark, restricting the scope to data from 2022 to 2024 and selecting only those associated with the 9 common error types listed in Table 3 with corresponding DQC US IDs. Leveraging the “sec-url” field in these reports, we then accessed the SEC (U.S. Securities and Exchange Commission) [^4] https://www.sec.gov/ database to obtain 218 corresponding XBRL filings, which constitute the foundation of our benchmark raw data.
@@ 4.1. Raw Data Collection
4.2. Data Annotation
Based on the filtered raw data described in Section 4.1, we annotated a total of 1,102 instances for the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark. As the first benchmark to focus on structured, multi-document, and taxonomy-driven inputs, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides a unique setting for evaluating LLMs. Table 4 shows that each task involves long and complex inputs, with average lengths exceeding 33k tokens per instance when measured using the “cl100k_base” GPT tokenizer. Among them, FinRE contains the largest total token volume and number of instances, while FinMR exhibits the longest average input length per instance. These statistics highlight the structural and relational complexity of XBRL filings, underscoring the need for careful annotation to ensure reliable evaluation.
Table 4. Statistics of input length (tokens) across tasks. Tokens are calculated using the “cl100k_base” tokenizer.
@@ 4.2. Data Annotation
4.2.1. Data Segmentation
For each complete XBRL filing, we extracted all statement IDs, which are their associated roleType IDs from the schema document, to segment the filing into distinct sections. Each section corresponds to a specific financial statement, such as the balance sheet or cash flow statement. Using these statement IDs, we then extracted the related content across linkbase documents. In particular, the presentation linkbase defines the logical and visual arrangement of financial elements within each statement and connects the facts in the instance document to their structured representation in reports. Following this process: from the schema document to the presentation linkbase, and subsequently to other linkbases and instance documents, we segmented each filing into multiple sections. These sections, treated as sub-filings of the original XBRL filing, preserve the inherent structure while reducing document length, thereby improving data manageability and enhancing the efficiency of LLM processing and evaluation.
4.2.2. Annotation with LLM Judgment
Each XBRL filing is retrieved using the ”sec-url” from the error messages, which provide explicit details, including the specific document components (e.g., calculation linkbase or instance document) where errors occur, the incorrect values, the invalid relationships, and the expected correct results. Therefore, we do the following processing for each task using the GPT-4o-mini model. We provide the specific prompts for each task in Appendix A.
(1) For the FinSM task, we extract all reported incorrect concepts from the error messages as the ground-truth answers and retrieve the corresponding sub-filing segments based on these concepts to serve as the input XBRL filing.
(2) For the FinRE task, we map the three defined relationship types to the corresponding error messages based on the DQC US ID. Specifically, DQC_0081 corresponds to “Reversal”, DQC_0001 to “CombinationErr”, and DQC_0145 to “Inappropriateness”. For each relationship type, we extract the participating elements from the error messages and construct relationship triples in the format: (head element, relation, tail element). The head and tail elements are used as inputs, while the relation serves as the ground-truth label. Similarly, based on the extracted elements, we retrieve the corresponding sub-filing segments to serve as the input XBRL filing.
(3) For the FinMR task, based on our task formulation, we annotate two key aspects for each question: the reported value in the instance document and the inferred generative value. Both answers are derived from the error message. Since the question in this task is concept-dependent, we extract the relevant concept and period from the error message to construct the input question. Similarly, we retrieve the corresponding sub-filing segments based on the extracted concepts to serve as the input XBRL filing.
4.2.3. US-GAAP Taxonomy Chunk
To facilitate structured access to the US-GAAP Taxonomy, we divided the taxonomy into concept-centric chunks. Each concept and its associated semantic information were extracted from the official taxonomy Excel files (e.g., GAAP Taxonomy 2024.xlsx). Specifically, we parsed all relevant sheets, including Concepts, Presentation, Calculation, Definition, Reference, and Enumeration, and organized them into two major chunk types: core chunks, which contain each concept’s intrinsic attributes such as label, type, balance, period type, and documentation, and relation chunks, which describe its structural and semantic links to other concepts. The relation chunks encode various dependency types, including hierarchical presentation paths, parent–child calculation relationships, definitional arcs, and external references. The details are shown in Appendix B. This chunking process preserves the hierarchical and relational structure of the taxonomy while enabling fine-grained indexing, retrieval, and reasoning at the concept level.
@@ 4.2.4. Quality Control
To ensure the reliability and consistency of the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, we conducted a series of quality control procedures across both data extraction and annotation phases. (1) A random sample of 50 filings was manually checked to verify that the extracted statement segments matched their roleType identifiers and preserved the original hierarchy. (2) For annotation, 50% of the instances were double-validated by comparing GPT-4o-mini outputs with independent human review, and discrepancies were corrected to ensure accuracy. (3) Finally, taxonomy chunks were cross-checked to ensure that every relation chunk referenced a valid core concept. These steps collectively guarantee structural fidelity and reliable ground-truth annotations for evaluating LLM reasoning over structured financial data.
4.3. Evaluation Metrics
Our goal is to explore the capabilities of current state-of-the-art LLMs on 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: and gain a deeper understanding of their strengths and limitations. To achieve this, we evaluate LLM performance across various aspects for each task. The details are shown in Appendix C.
For the FinSM task, we evaluate retrieval performance using Hit Rate (HR@kHR@k(He et al., 2017), Recall ($R@k$), and Macro-F1 ($MF1@k$), where k=1,5,10,20k={1,5,10,20}. Hit Rate measures whether the model retrieves at least one relevant item within the top-$k$ results. Recall assesses how completely the model retrieves all relevant elements. Macro-F1 balances precision and recall by averaging query-level F1 scores, providing an overall view of retrieval accuracy and coverage (Shani and Gunawardana, 2010).
For the FinRE task, we evaluate model performance using Accuracy (Acc), Macro Precision (P), Macro Recall (R), and Macro F1. Accuracy measures the overall proportion of correctly classified relations. Macro Precision and Macro Recall average performance across the three relation types. Macro F1 provides a balanced summary by combining precision and recall for each relation type and averaging the results.
For the FinMR task, we evaluated the mathematical reasoning capabilities of various LLMs using the LLM-as-a-judge approach (Gu et al., 2024; Li et al., 2024) to measure overall Accuracy. Additionally, we designed three error-rate indicators, including the structural error rate (SER), extraction error rate (EER), and calculation error rate (CER), to evaluate the quality of the generated answers by LLMs. Together, these metrics provide a comprehensive evaluation of mathematical reasoning performance.
4.4. Evaluation LLMs
Our goal is to evaluate the foundational capabilities of SOTA LLMs on the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, to better understand their strengths and limitations in financial auditing. To this end, we selected a diverse set of models: one closed-source general SOTA LLM (GPT-4o (Hurst et al., 2024)), ten open-source general SOTA LLMs of different sizes (DeepSeek-V3 (Liu et al., 2024), Qwen3-235B-A22B-Instruct-2507 (Team, 2025), Llama-4-Scout-17B-16E-Instruct (Meta, 2025), Qwen2.5-72B-Instruct (Yang et al., 2024), Llama-3.3-70B-Instruct (Grattafiori et al., 2024), Qwen3-32B (Team, 2025), gemma-3-27b-it (Team, 2024), gemma-3-12b-it (Team, 2024), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Llama-3.2-3B-Instruct (Grattafiori et al., 2024)), and two open-source financial SOTA LLMs (Fino1 (Qian et al., 2025) and FinR1 (Liu et al., 2025)). Detailed specifications of the evaluated models, including the exact versions used, are provided in Table 9 in the Appendix.
4.5. Evaluation Environments

### v2

@@ 3.1. Task Formulation
Given a target XBRL filing consisting of six interconnected, XML documents $\mathcal{D}=\{d_{1},d_{2},d_{3},d_{4},d_{5},d_{6}\}$ (described in Table 3) and a reference US-GAAP taxonomy $\mathcal{T}$, the task is to identify errors $\mathcal{E}$ within the filing. These errors include semantic-matching inconsistencies, relationship-extraction failures, and mathematical-reasoning violations. Models must analyze the filing’s internal structure and semantics while leveraging domain knowledge from the reference taxonomy. We formulate this as follows:
@@
where $\mathcal{E}$ is the specific description of the errors found in the $\mathcal{D}$.
To operationalize error detection across XBRL filings, we identify three fundamental capabilities: semantic consistency, structural relationship understanding, and numerical reasoning. Accordingly, 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: defines three tasks: (1) Financial Semantic Matching (FinSM) identifies mismatches between textual mentions in filings and standardized taxonomy labels; (2) Financial Relationship Extraction (FinRE) detects errors in hierarchical and compositional relationships among financial items; and (3) Financial Mathematical Reasoning (FinMR) verifies numerical correctness based on accounting logic and contextual inference.
3.1.1. FinSM
The semantic matching task evaluates whether models can identify financial tags in XBRL filings that misalign with standardized US-GAAP taxonomy concepts. Given a filing with multiple candidate tag assignments, the objective is to detect semantically inconsistent or inappropriate concept assignments within a set of complex, structured documents.
@@ 3.1.1. FinSM
This task follows an information retrieval paradigm. Given (1) a query $Q$ describing a financial term related to currency or credit risk concentration, (2) an XBRL filing $\mathcal{D}$, and (3) the US-GAAP taxonomy $\mathcal{T}$. The goal is to retrieve the set of mismatched US-GAAP tags within the filing, denoted as $\mathcal{E}$. For the FinSM task, we evaluate retrieval performance using metrics: Hit Rate (HR@kHR@k(He et al., 2017), Recall ($R@k$), and Macro-F1 ($MF1@k$), with $k\in\{1,5,10,20\}$.
3.1.2. FinRE
The relationship extraction task evaluates whether models can identify structural errors among financial elements by interpreting hierarchical and compositional dependencies between the filing and the external taxonomy. This task assesses the LLMs’ capacity for semantic understanding and structured data interpretation within the financial reporting context.
@@ 3.1.2. FinRE
Formally, given a relation error type space $L$ with predefined semantic definitions, an XBRL filing $\mathcal{D}$, and the US-GAAP taxonomy $\mathcal{T}$, the task aims to classify the specific relation error type $\ell\in L$ between two specific elements $e_{1}$ and $e_{2}$, which are derived from the current XBRL filing $\mathcal{D}$, based on the information provided by $\mathcal{T}$.
We define three relation error types $\{\ell_{1},\ell_{2},\ell_{3}\}$ in $L$: (1) Reversal, which occurs when the hierarchical relationship between parent and child elements is mistakenly reversed; (2) Inappropriateness, which indicates that a child element is incorrectly associated with an inappropriate parent element; and (3) CombinationErr, which refers to an invalid combination of axis and member elements that violates the structural constraints defined in the taxonomy. We use the metrics, including Accuracy (Acc), Macro Precision (P), Macro Recall (R), and Macro F1, to evaluate the performance of this task.
3.1.3. FinMR
The mathematical reasoning task focuses on inferring correct numerical values among financial elements based on structured representations, calculation hierarchies, and constraints defined in the XBRL filing. The objective is to assess whether reported values are internally consistent and compliant with domain-specific standards.
@@ 3.1.3. FinMR
Specifically, the task involves an XBRL filing $\mathcal{D}$, the US-GAAP taxonomy $\mathcal{T}$, and a pair of questions ($q_{1}$, $q_{2}$). Here, $q_{1}$ requests the extraction of a reported value and $q_{2}$ pertains to the computation of the corresponding true value. The model is required to: (1) extract the reported numeric value $v$ for a given instance from $\mathcal{D}$, and (2) compute the corresponding real value $\mu$ based on the structural and numerical information in $\mathcal{D}$ and $\mathcal{T}$. The extracted value $v$ and the computed value $\mu$ are then compared to determine whether the reported value is correct. The output of this task is a structured answer containing both $v$ and $\mu$, represented in a JSON format. For this task, we evaluate the overall Accuracy using an LLM-as-a-judge framework (Gu et al., 2024; Li et al., 2024), together with three complementary error indicators. The structural error rate (SER) measures cases where the model fails to produce a valid output structure in the required JSON format. The extraction error rate (EER) captures errors in identifying the correct numerical value to be extracted, even when the output structure is valid. The calculation error rate (CER) reflects mistakes in numerical computation, where the final calculated value is incorrect despite the correct structure and extraction.
3.2. Data Collection and Construction
0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: is established as a high-fidelity testbed to catalyze research on taxonomy-grounded reasoning and structured retrieval within complex, multi-document financial environments. As illustrated in Figure 2, our resource bridges the gap between raw XML data and the semantic constraints of the US-GAAP taxonomy.
3.2.1. Authoritative Error Signals from DQC Messages
To identify reliable auditing errors, we ground our benchmark construction in authoritative error signals released by the Data Quality Committee (DQC) [^5] https://xbrl.us/data-quality/filing-results/. Our construction follows a two-step pipeline: (1) Pre-screening: We collected 4,545 official DQC error messages for 372 companies between 2020 and 2024 via the XBRL US portal. Each message links to a deterministic DQC US rule, providing unambiguous labels for semantic, relational, or numerical inconsistencies. (2) Filtering: We selected the 9 most frequent DQC error types, covering 60.33% of all cases (Table 2), and retrieved 218 corresponding XBRL filings from the SEC EDGAR database, the official repository for public company disclosures.
@@ 3.2.1. Authoritative Error Signals from DQC Messages
3.2.2. XBRL Filing Segmentation
Using “SEC-url” identifiers from DQC messages (Section 3.2.1), we retrieve the corresponding XBRL filings as the primary data representation for structured financial reporting. An XBRL filing comprises multiple interconnected XML documents that jointly encode both financial facts and their structural constraints. As summarized in Table 3, each filing contains: an instance document with reported values, a schema document defining element types, and several linkbases (calculation, presentation, definition, and label) that specify hierarchical organization, arithmetic relationships, and semantic metadata (Debreceny et al., 2010; Hoitash and Hoitash, 2018).
Due to this rich but highly interdependent structure, complete XBRL filings are often extremely long and difficult to process directly, posing challenges for both manual inspection and long-context LLM evaluation. To balance document length with structural fidelity, we perform statement-level filing segmentation. Specifically, we extract roleType identifiers from the schema document to identify distinct financial statements (e.g., balance sheets or income statements). We then leverage the presentation linkbase to map the hierarchical layout of each statement to its corresponding facts in the instance document. This segmentation preserves the original XML dependencies across the schema, linkbases, and instance documents, while producing manageable, structure-aware sub-filings suitable for downstream auditing tasks and long-context LLM processing.
Table 3. Main Components of XBRL Filings
Component Description
3.2.3. US-GAAP Taxonomy Digitalization
To facilitate structured access to the US-GAAP Taxonomy, we digitalize it into concept-centric chunks. We extract each concept and its associated semantic information from the official taxonomy files (e.g., GAAP Taxonomy 2024.xlsx), parsing all relevant sheets: Concepts, Presentation, Calculation, Definition, Reference, and Enumeration. We organized the extracted information into two major chunk types: core chunks contain each concept’s intrinsic attributes, such as label, type, balance, period type, and documentation. relation chunks encode structural and semantic links to other concepts, including hierarchical presentation paths, parent–child calculation relationships, definitional arcs, and external references. This digitalization process preserves the taxonomy’s hierarchical and relational structure while enabling fine-grained concept-level indexing, retrieval, and reasoning.
3.3. Data Annotation and Quality Control
0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: provides 1,102 annotated instances, each averaging over 33k tokens (Table 4), capturing the long-context characteristics of practical financial auditing scenarios.
Table 4. Resource statistics across tasks (tokens via cl100k_base).
@@ 3.3. Data Annotation and Quality Control
3.3.1. Annotation via LLM-Assisted Parsing
Ground-truth supervision is deterministically derived from DQC error messages. We employ GPT-4o-mini solely as a high-efficiency parser, rather than a decision-maker, to convert these messages into structured annotations. The model extracts (i) DQC-defined ground-truth labels, (ii) task-specific elements for formulating queries, relations, or questions, and (iii) evidence pointers linking errors to relevant XBRL sub-filings and US-GAAP taxonomy chunks. These pointers enable the retrieval of the corresponding XBRL segments and taxonomy chunks, grounding each instance in both the reported filing and the governing rules. Task-specific instances are then constructed as follows: (1) FinSM: Semantically incorrect US-GAAP concepts identified in DQC messages serve as ground-truth answers. The corresponding statement-level XBRL sub-filings are retrieved as input. (2) FinRE: DQC rule IDs are deterministically mapped to relation labels (DQC_0081: Reversal, DQC_0001: CombinationErr, DQC_0145: Inappropriateness). Participating elements form $(\textit{head},\textit{relation},\textit{tail})$ triples, where the relation is the ground-truth label and the associated XBRL sub-filings provide input context. (3) FinMR: DQC messages contain both reported values and correct values implied by calculation linkbases. We use the target concept and reporting period to construct the input questions, with statement-level XBRL sub-filings serving as the input for numerical reasoning.
@@ 3.3.2. Quality Control
To ensure the reliability and consistency of the 0.07451 0.13725 0.6902F0.07059 0.19216 0.6549i0.06667 0.25098 0.61961n0.06275 0.30588 0.58431A0.05882 0.36471 0.54902u0.05882 0.41961 0.5098d0.0549 0.47843 0.47451i0.05098 0.53333 0.43922t0.04706 0.59216 0.40392i0.04314 0.64706 0.36863n0.03922 0.70588 0.33333g\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset:\__color_backend_reset: benchmark, we conducted a multi-stage quality control process: (1) Structural Verification: We manually inspected 50 randomly sampled filings to verify that extracted statement segments correctly matched their roleType identifiers and preserved the original reporting hierarchy. (2) Annotation Validation: 50% of all instances underwent double-validation, where GPT-4o-mini outputs were compared against independent human review, with all discrepancies corrected to ensure ground-truth accuracy. (3) Referential Integrity: We verified that all relation chunks reference valid core concepts in the taxonomy, ensuring the integrity of taxonomy-grounded reasoning paths.
3.4. Evaluated LLMs